### v1

@@ 6.1.2 Decay rate of the eigenfunctions
Suppose $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$ is a bounded domain with Lipschitz boundary and connected $\Omega^{\mathrm{ext}}$. Any harmonic function $u\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$ satisfies
@@
Suppose $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$ is a bounded domain with Lipschitz boundary and connected $\Omega^{\mathrm{ext}}$ and let $u_{1}\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$ be a eigenfunction associated with $\sigma_{1}\left(\Omega^{\mathrm{ext}}\right)$. Then,
$\lim_{|x|\to\infty}\frac{u_{1}(x)}{|x|^{2-n}}\neq 0.$
@@
If $\prod_{j=1}^{n-1}\kappa_{j}(s)=0$, the inequality becomes trivial. Thus, we assume $\kappa_{j}(s)>0$ for any $j\leq n-1$.
@@
To show that the infimum in (6.3) is attained, consider a minimising sequence $(f_{k})_{k\in\mathbb{N}}\subset H^{1}_{\mathrm{loc}}((0,\infty))$ with $f_{k}(0)=1$, $f_{k}(x)\to 0$ as $x\to\infty$ and $\lim_{k\to\infty}\int_{0}^{\infty}|f_{k}^{\prime}(t)|^{2}\zeta(s,t)\,\mathrm{d}t=K\left(s,\Omega^{\mathrm{ext}}\right)$. It is straightforward to show that, for each $R>0$, the sequence is bounded in $H^{1}((0,R))$. Hence, by standard compactness results, there exists a subsequence, again denoted by $f_{k}$, that converges weakly in $H^{1}((0,R))$ and strongly in $L^{2}((0,R))$ to a limit function $f\in H^{1}((0,R))$.
@@
where $c_{j}(s)=\frac{\kappa_{j}(s)^{n-2}}{\prod_{\begin{subarray}{c}i=1,\,i\neq j\end{subarray}}^{n-1}(\kappa_{j}(s)-\kappa_{i}(s))}$ are the coefficients of the partial fraction decomposition
$1=\sum_{j=1}^{n-1}c_{j}\prod_{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{n-1}(1+\kappa_{i}(s)r).$
To ensure that $f(t)$ is bounded, we require $\sum_{j=1}^{n-1}\frac{c_{j}}{\kappa_{j}(s)}=0$. Thus, the asymptotic behaviour of $f(t)$ simplifies to
$\lim_{t\to\infty}f(t)=C_{1}+C_{2}\log\left(\prod_{j=1}^{n-1}\kappa_{j}(s)^{\frac{c_{j}}{\kappa_{j}(s)}}\right),$
and the solution decays at infinity if and only if $C_{1}=-C_{2}\log\left(\prod_{j=1}^{n-1}\kappa_{j}(s)^{\frac{c_{j}}{\kappa_{j}(s)}}\right)$. Because $\sum_{j=1}^{n-1}c_{j}=1$, this yields
$K\left(s,\Omega^{\mathrm{ext}}\right)=\frac{-f^{\prime}(0)}{f(0)}=\frac{1}{\log\left(\prod_{j=1}^{n-1}\kappa_{j}(s)^{\frac{c_{j}}{\kappa_{j}(s)}}\right)}=\frac{1}{\sum_{j=1}^{n-1}\frac{c_{j}}{\kappa_{j}(s)}\log\left(\kappa_{j}(s)\right)}.$
Substituting the expression for $c_{j}$, we arrive at
$K\left(s,\Omega^{\mathrm{ext}}\right)=\frac{1}{\sum_{j=1}^{n-1}\frac{\kappa_{j}(s)^{n-3}}{\prod_{\begin{subarray}{c}i=1,\,i\neq j\end{subarray}}^{n-1}(\kappa_{j}(s)-\kappa_{i}(s))}\log\left(\kappa_{j}(s)\right)},$
which completes the proof when the curvatures $\kappa_{j}(s)$ are distinct. When they are not, we simply proceed by approximation. ∎
@@
For distinct, positive $\alpha_{1},\ldots,\alpha_{n-1}$, by [Pi85, Proposition 4.1, Theorem 3.8],
$L(\alpha_{1},\ldots,\alpha_{2})\geq(n-2)\sqrt[n-1]{\prod_{j=1}^{n-1}|\alpha_{j}(s)|}.$
@@
The first inequality in (6.4) can be an equality only if $u$ is an eigenfunction; since the first eigenfunction does not change sign (see §6.1.1), the second inequality can only be equal if $K\left(s,\Omega^{\mathrm{ext}}\right)$ is constant. By (1.4) and (6.6), the geometric mean of the curvatures must also be constant. By [Mu87, Theorem 2], this forces $\partial\Omega$ to be a sphere. ∎
@@
6.3.1 Proof of Theorem 1.16
@@
where the function $\omega:\partial B_{1}\to\mathbb{R}$ is defined by
@@
In case of equality, arguing in the same was as in [GirPo10b, Section 2.3], we get that each $x_{i}$ is an eigenfunction of (6.13) with eigenvalue one, and it follows that $\omega(z)=1$ for all $z\in\partial B_{1}$. Let $a:=\psi^{-1}(0)\in B_{1}$, and consider the conformal automorphism $\delta_{a}:B_{1}\to B_{1}$ of the unit disk defined by
@@
Because $\psi$ is a biholomorphic map, $\psi^{\prime}(z)$ does not vanish for $z\in B_{1}$. Moreover, $a$ is the unique zero of $\psi$, and it is simple. It follows that $\frac{1}{\psi(z)}$ is a meromorphic function with a simple pole at $a$. Therefore, the function $\upsilon(z):=\frac{\psi^{\prime}(z)}{\psi(z)^{2}}\delta_{a}(z)^{2}$ has a removable singularity at $a$, and hence it can be extended to a holomorphic function on $B_{1}$. Moreover, $|\upsilon(z)|=1$ for each $z\in\partial B_{1}$. By the maximum modulus principle, this implies that there is a constant $c$ with $|c|=1$, and such that $\upsilon(z)=c$ for all $z\in B_{1}$. As a consequence, for any $z\in B_{1}\setminus\{a\}$ we have,
@@
and $|C|<1$ since $\psi$ is holomorphic. It is now easy to see that $\Omega^{*}=\psi(B_{1})$ is a disk, and hence $\Omega$ is a disk as well. This completes the proof of the theorem. ∎
@@ 6.3.2 Higher dimensions
$\min_{x\in\mathscr{e}_{a}^{(2)}}L(\kappa_{1}(x),\ldots,\kappa_{n-1}(x))=\frac{1}{-\log(a)\frac{a}{(1-a^{2})^{n-2}}-\frac{a}{(n-3)!}\frac{\mathrm{d}^{n-3}}{\mathrm{d}t^{n-3}}\left(\frac{t^{n-3}\log(t)}{1-at}\right)\Big|_{t=a}}.$
@@
$\begin{split}\min_{x\in\mathscr{e}_{a}^{(2)}}L(\kappa_{1}(x),\ldots,\kappa_{n-1}(x))&=\frac{1}{-\log(a)\frac{a}{(1-a^{2})^{n-2}}-\frac{a}{(1-at)^{n-2}}\log(a)+O(a)}\\ &=\frac{(1-a^{2})^{n-2}}{-2a\log(a)+O(a)}\qquad\text{as }a\searrow 0.\end{split}$
@@
For $k>2$, the argument carries over in the same way, leading again to $\sigma_{1}\left((\tilde{\mathcal{E}}_{a}^{(k)})^{\mathrm{ext}}\right)\to\infty$ as $a\searrow 0$. In fact, the divergence is faster, which can be understood informally from the fact that the principal curvature $\frac{1}{a\sqrt{D(x)}}$ is weighted more heavily in the logarithmic mean as $k$ increases. ∎

### v2

@@ 6.1.2 Decay rate of the eigenfunctions
Suppose $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$, is a bounded domain with Lipschitz boundary and connected $\Omega^{\mathrm{ext}}$. Any harmonic function $u\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$ satisfies
@@
Suppose $\Omega\subset\mathbb{R}^{n}$, $n\geq 3$, is a bounded domain with Lipschitz boundary and connected $\Omega^{\mathrm{ext}}$, and let $u_{1}\in E^{1}\left(\Omega^{\mathrm{ext}}\right)$ be a eigenfunction associated with $\sigma_{1}\left(\Omega^{\mathrm{ext}}\right)$. Then, there exists a $c\in\mathbb{R}\setminus\{0\}$ such that
$\lim_{|x|\to\infty}\frac{u_{1}(x)}{|x|^{2-n}}=c.$
@@
If $\prod_{j=1}^{n-1}\kappa_{j}(s)=0$, the inequality becomes trivial because the logarithmic mean vanishes when one of the curvatures vanishes. Thus, we assume $\kappa_{j}(s)>0$ for any $j\leq n-1$.
@@
To show that the infimum in (6.3) is attained, consider a minimising sequence $(f_{k})_{k\in\mathbb{N}}\subset H^{1}_{\mathrm{loc}}((0,\infty))$ with $f_{k}(0)=1$, $f_{k}(x)\to 0$ as $x\to\infty$ and $\lim_{k\to\infty}\int_{0}^{\infty}|f_{k}^{\prime}(t)|^{2}\zeta(s,t)\,\mathrm{d}t=K\left(s,\Omega^{\mathrm{ext}}\right)$. It is straightforward to show that, for each $R>0$, the sequence is bounded in $H^{1}((0,R))$. Hence, by standard compactness results, there exists a subsequence, again denoted by $f_{k}$, that converges weakly in $H^{1}((0,R))$ and strongly in $L^{2}((0,R))$ to a limit function $f\in H^{1}((0,R))$. Specifically, $f(0)=1$.
@@
where
$c_{j}(s)=\frac{\kappa_{j}(s)^{n-2}}{\prod_{\begin{subarray}{c}i=1,\,i\neq j\end{subarray}}^{n-1}(\kappa_{j}(s)-\kappa_{i}(s))},\qquad j=1,\dots,n-1,$ (6.6)
are the coefficients of the partial fraction decomposition
$1=\sum_{j=1}^{n-1}c_{j}(s)\prod_{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{n-1}(1+\kappa_{i}(s)r).$
It is easy to check that $\sum_{j=1}^{n-1}\frac{c_{j}(s)}{\kappa_{j}(s)}=0$, which ensures that $f(t)$ is bounded. Thus, the asymptotic behaviour of $f(t)$ simplifies to
$\lim_{t\to\infty}f(t)=C_{1}+C_{2}\log\left(\prod_{j=1}^{n-1}\kappa_{j}(s)^{\frac{c_{j}(s)}{\kappa_{j}(s)}}\right),$
and the solution decays at infinity if and only if $C_{1}=-C_{2}\log\left(\prod_{j=1}^{n-1}\kappa_{j}(s)^{\frac{c_{j}(s)}{\kappa_{j}(s)}}\right)$. Because $\sum_{j=1}^{n-1}c_{j}(s)=1$, this yields
$K\left(s,\Omega^{\mathrm{ext}}\right)=\frac{-f^{\prime}(0)}{f(0)}=\frac{1}{\log\left(\prod_{j=1}^{n-1}\kappa_{j}(s)^{\frac{c_{j}(s)}{\kappa_{j}(s)}}\right)}=\frac{1}{\sum_{j=1}^{n-1}\frac{c_{j}(s)}{\kappa_{j}(s)}\log\left(\kappa_{j}(s)\right)}.$
Substituting the expressions (6.6) for $c_{j}(s)$, and taking into account (1.3), we arrive at
$K\left(s,\Omega^{\mathrm{ext}}\right)=\frac{1}{\sum_{j=1}^{n-1}\frac{\kappa_{j}(s)^{n-3}}{\prod_{\begin{subarray}{c}i=1,\,i\neq j\end{subarray}}^{n-1}(\kappa_{j}(s)-\kappa_{i}(s))}\log\left(\kappa_{j}(s)\right)}=(n-2)L\left(\kappa_{1}(s),\dots,\kappa_{n-1}(s)\right),$
which completes the proof when the curvatures $\kappa_{j}(s)$ are distinct. As our formula for $K(s,\Omega^{\mathrm{ext}})$ extends continuously to the case when curvatures coincide, the inequality remains valid in that case as well. ∎
@@
For distinct, positive $\alpha_{1},\ldots,\alpha_{n-1}$, by [Pi85, Proposition 4.1 and Theorem 3.8],
$L(\alpha_{1},\ldots,\alpha_{2})\geq\sqrt[n-1]{\prod_{j=1}^{n-1}|\alpha_{j}(s)|}.$
@@
The first inequality in (6.4) can be an equality only if $u$ is an eigenfunction. Since the first eigenfunction does not change sign (see §6.1.1), the second inequality in (6.4) can only be an equality if $K\left(s,\Omega^{\mathrm{ext}}\right)$ is constant. By (1.5) and (6.7), then the geometric mean of the curvatures must also be a constant. By [Mu87, Theorem 2], this forces $\partial\Omega$ to be a sphere. ∎
@@
6.3.1 Proof of the isoperimetric inequality in dimension two
Proof of Theorem 1.16.
@@ 6.3.1 Proof of the isoperimetric inequality in dimension two
where the function $\omega:\partial B_{1}\to\mathbb{R}$ is given by
@@
In case of equality, arguing in the same way as in [GirPo10b, Section 2.3] and [FrLa20, pp. 1038-1039], we get that each $x_{i}$ is an eigenfunction of the weak version of (6.14) corresponding to eigenvalue one, and it follows that $\omega(z)=1$ for almost all $z\in\partial B_{1}$. Let $a:=\psi^{-1}(0)\in B_{1}$, and consider the conformal automorphism $\delta_{a}:B_{1}\to B_{1}$ of the unit disk defined by
@@
Because $\psi$ is a biholomorphic map, $\psi^{\prime}(z)$ does not vanish for $z\in B_{1}$. Moreover, $a$ is the unique zero of $\psi$, and it is simple. It follows that $\frac{1}{\psi(z)}$ is a meromorphic function with a simple pole at $a$. Therefore, the function $\upsilon(z):=\frac{\psi^{\prime}(z)}{\psi(z)^{2}}\delta_{a}(z)^{2}$ has a removable singularity at $a$, and hence it can be extended to a holomorphic function on $B_{1}$. Note that $\upsilon$ does not vanish in $B_{1}$. Moreover, $|\upsilon(z)|=1$ for almost all $z\in\partial B_{1}$. Hence, the function $z\mapsto\log(|\upsilon(z)|)$ is harmonic on $B_{1}$ and it vanishes almost everywhere on $\partial B_{1}$. Since $\Omega$ is Lipschitz, the conformal map $\psi$ is regular enough for $\log|\upsilon|$ to admit a representation as a Poisson integral on $\partial B_{1}$, hence $\log|\upsilon|$ vanishes everywhere on $B_{1}$, see [FrLa20]. This implies the existence of a constant $c$ with $|c|=1$, such that $\upsilon(z)=c$ for all $z\in B_{1}$. As a consequence, for any $z\in B_{1}\setminus\{a\}$ we have,
@@
and $|C|<1$ since $\psi$ is holomorphic. It is now easy to see that $\Omega^{*}=\psi(B_{1})$ is a disk, and hence $\Omega$ is a disk as well. ∎
@@ 6.3.2 Higher dimensions
$\min_{x\in\mathscr{e}_{a}^{(2)}}L(\kappa_{1}(x),\ldots,\kappa_{n-1}(x))=\frac{1}{(n-2)\left(-\log(a)\frac{a}{(1-a^{2})^{n-2}}-\frac{a}{(n-3)!}\left.\frac{\mathrm{d}^{n-3}}{\mathrm{d}t^{n-3}}\left(\frac{t^{n-3}\log(t)}{1-at}\right)\right|_{t=a}\right)}.$
@@
$\min_{x\in\mathscr{e}_{a}^{(2)}}L(\kappa_{1}(x),\ldots,\kappa_{n-1}(x))=\frac{(1-a^{2})^{n-2}}{(n-2)\left(-2a\log(a)+O(a)\right)}\qquad\text{as }a\searrow 0.$
@@
For $k>2$, the argument carries over in the same way, leading again to $\sigma_{1}\left(\left(\tilde{\mathcal{E}}_{a}^{(k)}\right)^{\mathrm{ext}}\right)\to\infty$ as $a\searrow 0$. In fact, the divergence is faster, which can be understood informally from the fact that the principal curvature $\frac{1}{a\sqrt{D(x)}}$ is weighted more heavily in the logarithmic mean as $k$ increases. ∎